# WAFFLED: Exploiting Parsing Discrepancies to Bypass Web Application Firewalls

### v2

@@ Abstract.
Web Security, Web Application Firewalls, Fuzzing, HTTP, Parsing Discrepancies, Request Smuggling, WAF Evasion
* Corresponding Author
1. Introduction
@@
As the threat landscape evolves, HTTP Request Smuggling (HRS) attacks have gained significant attention. HRS exploits discrepancies in the interpretation of HTTP requests between different entities in the communication chain, such as servers, proxies, and WAFs. These attacks can have severe consequences, including unauthorized access to sensitive information, session hijacking, and server compromise. The increasing complexity of web applications and their reliance on intermediary components have increased the risk of HRS vulnerabilities. Recent studies have investigated the importance of HRS, identifying new variants of attack and suggesting defense mechanisms (Grenfeldt et al., 2021; Kettle, 2020; Klein, 2020a, b).
@@
Availability. WAFFLED, fuzzing grammars, WAF configuration, web application frameworks and parsers code, and HTTP-Normalizer are open source, and are available on this  [^1] https://github.com/sa-akhavani/waffled/ repository.
2. Background
2.1. Web Content Types and RFCs Explained
@@
Our work focuses on using specific features of content-types to bypass a malicious request; thus, it is crucial to inspect the details of the content and media types discussed in this paper.
@@
RFCs (Request for Comments). RFCs (Internet Engineering Task Force (2025), IETF) are a series of documents that define protocols, procedures, and conventions used on the Internet and networking standards. They are published by the Internet Engineering Task Force (IETF) and related organizations. Each RFC is assigned a unique number, and these documents serve as the authoritative source of information on various Internet standards and protocols. For instance, the structure of HTTP headers, MIME types, and various other web technologies are defined by respective RFCs.
@@
ABNF (Augmented Backus-Naur Form). ABNF grammar rules (Crocker and Overell, 2008) are a formal notation used to specify the syntax of RFCs. ABNF extends the basic BNF notation to provide a more flexible and precise way to define the syntax of Internet protocols. It is particularly important in our work as it allows for precise definitions of content types and headers, which we exploit to identify parsing discrepancies in WAFs. By understanding and manipulating ABNF rules, we can generate requests that challenge the ability of WAFs to consistently parse and enforce these standards.
@@
Multipart. Multipart content-types allow HTTP messages to include multiple entities with varying media types in a single message body. They are crucial for handling file uploads and complex data structure support. These content-types, such as multipart/form-data, multipart/related, and multipart/mixed, are governed by specific RFCs (e.g., RFC 2387, RFC 7578, RFC 2045-2047) (Masinter, 2015; Levinson, 1998; Freed and Borenstein, 1996b, a; Moore, 1996) that define their structure and usage scenarios. Our research investigates the complexities of multipart/form-data requests, exploiting features such as boundary definitions, charset, content-disposition, and other custom headers.
@@
XML. Extensible Markup Language serves as a flexible format for structured data representation, commonly used in data exchange protocols. XML-based requests, defined in RFC 7303 (Thompson and Lilley, 2014), are characterized by elements such as DOCTYPE declarations, schemas, and CDATA sections. Our approach includes mutating these structural elements to investigate WAF responses to XML-specific parsing matters. For example, variations in DOCTYPE declarations and CDATA usage are tested to identify vulnerabilities in WAF handling of XML content.
@@
JSON. JavaScript Object Notation, described in RFC 8259 (Bray, 2017), is a lightweight data interchange format widely used in modern web applications. JSON requests are characterized by their simple structure based on key-value pairs. We manipulate JSON object formatting and nested structures to assess WAF handling of JSON parsing anomalies.
2.2. How WAFs Work
To understand the functioning of WAFs, one must visualize the transactional exchange between web applications and users. Before a web application receives an HTTP request, the WAF intercepts this request, serving as a firewall positioned between the application and the user. It thoroughly scrutinizes the request by analyzing its parameters, headers, and payload to identify any potential signs of malicious intent. If the request appears legitimate, the WAF allows it to pass through to the web application, enabling normal operation. However, if the request exhibits characteristics of an attack, the WAF either completely blocks the whole request, or removes the malicious part of it, preventing the malicious code from executing and protecting the web application from harm. Figure 1 represents this process. There are many industrial and open source WAFs available to the public, each using their own techniques to prevent malicious requests to be passed to the protected web application. Notable WAF providers include Cloudflare, Google Cloud Armor, Microsoft Azure WAF, Amazon AWS WAF, and ModSecurity.
Figure 1. WAF filtering malicious requests.
2.3. Foundations of WAF Operations and Ruleset Definitions
A WAF operates by inspecting HTTP traffic between a client and a web application, filtering out harmful requests to prevent attacks such as SQL injection, cross-site scripting (XSS), and malicious script uploads. The WAF is positioned inline with the traffic flow, allowing it to analyze and intercept malicious content before it reaches the application server. WAFs employ several detection approaches.
@@ 2.3. Foundations of WAF Operations and Ruleset Definitions
Managed Rulesets. WAFs use predefined sets of rules that define patterns associated with common attacks. These rulesets are frequently updated to respond to new threats. Most WAFs use their own custom rulesets. But one of the famous rulesets that is widely used among almost all WAFs and most custom rulesets use it as a base, is OWASP CRS (Core Rule Set) (owa, 2024). This ruleset is a set of generic attack detection rules for use with compatible WAFs. It aims to protect web applications from a wide range of attacks, including the OWASP Top Ten. CRS provides protection against many common attack categories, including SQL Injection, Cross Site Scripting, Local File Inclusion, etc.
@@
Machine Learning-Based Detections. Some advanced WAFs leverage machine learning algorithms to detect anomalies in traffic patterns. These models can adapt to new threats over time, offering a dynamic defense mechanism.
A WAF operates by scanning incoming HTTP requests, analyzing headers, cookies, URL parameters, and body content. Each request is parsed to extract these fields, allowing the WAF to evaluate them against a predefined set of rules. These rules follow a structured format:
@@
Listing 1: WAF Rule Format
@@
To accommodate complex attack patterns, WAF rules often rely on regular expressions (regex). For instance, detecting SQL injection attempts may involve scanning URL parameters or request bodies for SQL-related keywords. When a request arrives, the WAF extracts relevant fields, evaluates them against the defined rules, and executes the appropriate action. If a match is found, the WAF may block the request, log the event, or challenge the user with additional security measures such as a CAPTCHA. In cases where no match occurs, the request is forwarded to the application without interference.
Listing 2 presents an example of how a WAF rule might be defined and applied. First, there should be a rule definition.
@@
Listing 2: Example WAF Rule. Field: url_parameter user_input, Operator: contains. Value: DROP TABLE, and Action: block
After a request is received by a WAF, it does request Parsing and Rule Application. If a request is received with the URL parameter: user_input=DROP TABLE users, the WAF extracts the user_input parameter, and evaluates it against the rule. if "DROP TABLE" in user_input: The regex ”DROP TABLE” matches the parameter value. Since the action for such pattern is defined as ”block”, the WAF executes the block action, denying the request and logging the incident.
By using well-defined rules and regular expressions, a WAF effectively protects web applications from a wide range of attacks. The flexibility in defining rules allows security teams to tailor the WAF to the specific needs of their applications, enhancing the security without significantly impacting performance.
As explained, a critical aspect of WAF operation is how it parses and interprets incoming data. Parsing behaviors can differ significantly across various WAFs, leading to discrepancies in how they process the same set of data. These discrepancies present a potential avenue for bypassing WAFs, as attackers can exploit differences in parsing to craft requests that are interpreted harmlessly by the WAF, but maliciously by the target application. This work primarily focuses on uncovering new bypassing techniques by abusing these parsing discrepancies.
3. Related Work
3.1. Architecture, Payload, and Protocol Level Bypasses.
@@ 3. Related Work
At the architectural level, bypasses exploit vulnerabilities within the web application’s infrastructure. Techniques such as IP spoofing and server-side request forgery (SSRF) allow attackers to access the origin server directly, circumventing WAF protection. Cache poisoning attacks, as discussed by Meiners et al. (Nguyen et al., 2019), exploit CDN vulnerabilities to disrupt access to web applications. Additionally, file processing exploits, such as those highlighted by Freiling et al. (Jana and Shmatikov, 2012), offer insights into similar architectural vulnerabilities.
@@
Bypassing WAFs at the protocol level typically involves manipulating communication protocols. HTTP request smuggling, where a malicious request is embedded within a legitimate one, is a key example, with T-Reqs (Jabiyev et al., 2021) providing methods to identify such vulnerabilities. Similarly, features such as chunked transfer encoding, which involves fragmenting HTTP data into smaller chunks, can be exploited to evade WAF detection. Dalili’s work on IP fragmentation (sor, 2021) explore these protocol-level evasion strategies. Also, SYMTCP (Wang et al., 2020) focuses on evading deep packet inspection, a principle that can be applied to circumvent WAFs at the protocol level. Furthermore, Handley et al. (Handley et al., 2001) inspects the evasion of network intrusion detection systems, a concept closely related to bypassing WAFs at the protocol level.
Figure 2. WAFFLED Overview.
@@
Payload level bypassing techniques involve transforming the original payload to render it undetectable by the WAF. Obfuscation techniques, such as encoding or encryption, are frequently employed to conceal malicious code from WAF scrutiny. Additionally, evasion techniques, such as altering the order of characters or incorporating whitespace, can be used to confuse the WAF, and facilitate bypassing. Payload size manipulation, such as oversized POST requests that exploited a vulnerability in Google Cloud Armor (noa, 2022), illustrates how payload modifications can defeat WAF protections.
Not all bypasses are feasible using by targeting a single level approach above. Some works involve combining multiple techniques to achieve a bypass. AutoSpear (Qu et al., [n. d.]) exemplifies a tool that automates WAF bypassing using a combination of architectural, protocol, and payload-level techniques. Our work also fits in this category, since we use both features of protocol-level and payload-level techniques, although not modifying the attack payload.
@@
3.2. Differences with Prior Work on WAF Bypasses
In recent work, Wang et al. (Wang et al., 2024), presented WAFManis, in which they examined bypassing WAFs using protocol-level evasion. While this work aligns with our work in exploring WAF vulnerabilities, our research fundamentally differs in several key aspects.
Scope. While previous work has primarily concentrated on protocol-level attacks, our work uniquely targets content-type-specific attacks by exploiting parsing discrepancies within HTTP requests. That is, our approach extends beyond the protocol layer to focus on inconsistencies in both headers and bodies, targeting a wider array of bypass scenarios. Although there is some overlap in our findings with Wang et al., particularly concerning request header mutations, this is not the primary focus of our research. In contrast, by systematically examining all web-used content-types, our approach offers a broader scope, potentially identifying bypass techniques that have not been previously explored. Both works advocate for a systematic approach, yet our focus on content-type-specific parsing discrepancies enables the identification of a novel category of WAF bypasses.
Methodological Differences and Expanded Coverage. Although Wang et al. employ fuzzing to uncover WAF bypasses, the methodologies diverge significantly. Our approach emphasizes content-type-specific RFC rules and features within request bodies, such as application/xml namespaces, multipart/form-data custom headers and charsets, and application/json structures. This distinction introduces a systematic exploration of bypasses through various content-types, leading to the discovery of new categories of bypasses beyond what is covered in previous work.
Classification. We propose a classification system for bypass techniques, dividing them into 24 distinct types based on content types, which facilitates easier detection, analysis, and reporting. In contrast, Wang et al. defines only 3 attack types, which lack the granularity needed for in-depth study and mitigation. Our system also enables the identification of potential defenses for each bypass, as discussed in the evaluation.
Practicality. Our research demonstrates the real-world applicability of our findings by testing bypass techniques on 100 live websites, while Wang et al.’s work was limited to a controlled environment, highlighting the broader relevance of our work.
Mitigation. We introduce the HTTP-Normalizer, a proof-of-concept strategy designed to eliminate parsing discrepancies by enforcing strict RFC compliance, a contribution not present in previous work.
4. WAFFLED Design and Methodology
Figure 2 provides an overview of the WAFFLED approach for finding discrepancy-based bypass vectors for WAFs. The methodology involves generating, mutating, prefiltering, and testing HTTP requests across the WAFs and web application frameworks to identify successful bypass instances.
4.1. Input Generation and Mutation
The fuzzer generates valid HTTP requests using a predefined grammar that contains a web attack payload (e.g., SQL injection, Cross-Site Scripting). Mutations are applied everywhere except the attack payload itself, ensuring that bypass attempts focus on discrepancies rather than payload obfuscation. The highlighted part in Listing 3 shows an XSS attack payload, which remains unmodified during mutations.
These mutations are expected to give us new discrepancy-based bypass instances. The goal is to confuse WAFs so they fail to properly parse the mutated requests and allow them through, while the web application frameworks, despite the mutations, correctly interpret and execute the attack payload.
4.2. Identifying Successful Bypass Instances
@@ 4.3. Motivating Example
Listing 3: A malicious HTTP request using multipart/form-data content type with an XSS attack payload. The malicious part of the payload is highlighted and unmodified during the mutations.
@@
5boundary=fake-boundary;boundary*0=real-;boundary*1=boundary
6
7--fake-boundary
8Content-Disposition: form-data; name="field1"
9
10value1
11--fake-boundary--
12--real-boundary
13Content-Disposition: form-data: name="id"
14
15<script>alert(document.cookie)</script>
16--real-boundary--
@@
5. Experiment Setup
5.1. Tools and Infrastructure for Experiments
We used a modified version of the T-Reqs (Jabiyev et al., 2021) fuzzer for the experiments. The modifications include adding support for encrypted HTTP requests, and setting content-length dynamically based on the length of the body. We also modified the T-Reqs source code to avoid using < and > characters to parse the grammar, since these characters were also used in the XSS attack payload that was present in generated requests.
@@ 5.2. WAF Configuration and Setup
Table 1. List of Tested WAFs and Their Supported Request Content-Types (As of July 2024).
@@
Table 1 lists the WAFs and their supported content-types in our study.
@@
5.3. Web Application Frameworks Setup

### v3

@@ Abstract
1 Introduction
@@
As the threat landscape evolves, HTTP Request Smuggling (HRS) attacks have gained significant attention. HRS exploits discrepancies in the interpretation of HTTP requests between different entities in the communication chain, such as servers, proxies, and WAFs. These attacks can have severe consequences, including unauthorized access to sensitive information, session hijacking, and server compromise. The increasing complexity of web applications and their reliance on intermediary components have increased the risk of HRS vulnerabilities. Recent studies have investigated the importance of HRS, identifying new variants of attack and suggesting defense mechanisms [1, 2, 3, 4].
@@
Availability. Our results and source code, including fuzzer input grammars and WAF configurations are publicly available [^1] https://github.com/sa-akhavani/waffled. Sensitive bypass requests will remain restricted until the vulnerabilities are resolved.
2 Background
2.1 Web Content Types and RFCs Explained
@@
Our work focuses on using specific features of content-types to bypass a malicious request. Thus, it is crucial to inspect the details of the content and media types discussed in this paper.
@@
RFCs (Request for Comments). RFCs [5] are a series of documents that define protocols, procedures, and conventions used on the Internet and networking standards. They are published by the Internet Engineering Task Force (IETF) and related organizations. Each RFC is assigned a unique number, and these documents serve as the authoritative source of information on various Internet standards and protocols. For instance, the structure of HTTP headers, MIME types, and various other web technologies are defined by respective RFCs.
@@
ABNF (Augmented Backus-Naur Form). ABNF grammar rules [6] are a formal notation used to specify the syntax of RFCs. ABNF extends the basic BNF notation to provide a more flexible and precise way to define the syntax of Internet protocols. It is particularly important in our work as it allows for precise definitions of content types and headers, which we exploit to identify parsing discrepancies in WAFs. By understanding and manipulating ABNF rules, we can generate requests that challenge the ability of WAFs to consistently parse and enforce these standards.
@@
Multipart. Multipart content-types allow HTTP messages to include multiple entities with varying media types in a single message body. They are crucial for handling file uploads and complex data structure support. These content-types, such as multipart/form-data, multipart/related, and multipart/mixed, are governed by specific RFCs (e.g., RFC 2387, RFC 7578, RFC 2045-2047) [7, 8, 9, 10, 11] that define their structure and usage scenarios. Our research investigates the complexities of multipart/form-data requests, exploiting features such as boundary definitions, charset, content-disposition, and other custom headers.
@@
XML. Extensible Markup Language serves as a flexible format for structured data representation, commonly used in data exchange protocols. XML-based requests, defined in RFC 7303 [12], are characterized by elements such as DOCTYPE declarations, schemas, and CDATA sections. Our approach includes mutating these structural elements to investigate WAF responses to XML-specific parsing matters. For example, variations in DOCTYPE declarations and CDATA usage are tested to identify vulnerabilities in WAF handling of XML content.
@@
JSON. JavaScript Object Notation, described in RFC 8259 [13], is a lightweight data interchange format widely used in modern web applications. JSON requests are characterized by their simple structure based on key-value pairs. We manipulate JSON object formatting and nested structures to assess WAF handling of JSON parsing anomalies.
How WAFs Work. A WAF operates by inspecting HTTP traffic between clients and web applications, filtering out potentially harmful requests to prevent attacks such as SQL injection, cross-site scripting (XSS), and malicious script uploads. Positioned inline with the traffic flow, the WAF acts as a barrier between users and the application server, and is able to analyze requests in real-time before they reach the protected application. To make filtering decisions, WAFs operate by scanning incoming HTTP requests, analyzing headers, cookies, URL parameters, and body content. Figure 1 illustrates this process. Numerous commercial and open-source WAFs implement these mechanisms, including Cloudflare, Google Cloud Armor, Microsoft Azure WAF, Amazon AWS WAF, and ModSecurity. WAFs typically rely on one or more detection strategies:
Figure 1: WAF filtering malicious requests.
Figure 2: WAFFLED Overview.
@@
Managed Rulesets. WAFs use predefined sets of rules that define patterns associated with common attacks. These rulesets are frequently updated to respond to new threats. Most WAFs use their own custom rulesets. But one of the famous rulesets that is widely used among almost all WAFs and most custom rulesets use it as a base, is OWASP CRS (Core Rule Set) [14]. This ruleset is a set of generic attack detection rules for use with compatible WAFs. It aims to protect web applications from a wide range of attacks, including the OWASP Top 10. CRS provides protection against many common attack categories, including SQL Injection, Cross Site Scripting, Local File Inclusion, etc.
@@
Machine Learning-Based Detections. Advanced WAFs incorporate Machine Learning models that analyze traffic behavior and adapt over time, improving their ability to detect previously unseen threats.
To accommodate complex attack patterns, WAF rules often rely on regular expressions (regex). For instance, detecting SQL injection attempts may involve scanning URL parameters or request bodies for SQL-related keywords. When a request arrives, the WAF extracts relevant fields, evaluates them against the defined rules, and executes the appropriate action. If a match is found, the WAF may block the request, log the event, or challenge the user with additional security measures such as a CAPTCHA. In cases where no match occurs, the request is forwarded to the application without interference. By using well-defined rules and regular expressions, a WAF effectively protects web applications from a wide range of attacks.
Each request is parsed to extract these fields, allowing the WAF to evaluate them against a predefined set of rules. These rules follow a structured format that are displayed in listing 1 and are as follows:
@@
Listing 1: WAF Rule Format.
@@
Example of a WAF in action. Listing 2 illustrates a simple WAF rule definition and how it is applied during request inspection.
@@
Listing 2: Example WAF Rule.
In this case, when a request with the URL parameter user_input=DROP TABLE users arrives, the WAF parses the request and extracts relevant fields. Then, it evaluates the value of user_input against the defined rule. The pattern "DROP TABLE" is matched using a regular expression, triggering the block action. As a result, the WAF denies the request and logs the incident for further analysis.
While such rule-based filtering mechanisms are effective in many scenarios, a critical but often overlooked aspect of WAF behavior is how the system parses and interprets incoming data. Different WAFs may handle HTTP request parsing in different ways, particularly when dealing with outdated RFCs, nested data structures, or uncommon content types. These parsing discrepancies, create opportunities for attackers to bypass protections by crafting inputs that are marked as valid by the WAF but would exploit the the target application. This work focuses on uncovering and categorizing such evasions by analyzing how parsing discrepancies arise and demonstrating practical techniques that abuse these mismatches to bypass modern WAFs.
3 Related Work
@@
At the architectural level, bypasses exploit vulnerabilities within the web application’s infrastructure. Techniques such as IP spoofing and server-side request forgery (SSRF) allow attackers to access the origin server directly, circumventing WAF protection. Cache poisoning attacks, as discussed by Meiners et al. [15], exploit CDN vulnerabilities to disrupt access to web applications. Additionally, file processing exploits, such as those highlighted by Freiling et al. [16], offer insights into similar architectural vulnerabilities.
@@
Bypassing WAFs at the protocol level typically involves manipulating communication protocols. HTTP request smuggling, where a malicious request is embedded within a legitimate one, is a key example, with T-Reqs [17] providing methods to identify such vulnerabilities. Similarly, features such as chunked transfer encoding, which involves fragmenting HTTP data into smaller chunks, can be exploited to evade WAF detection. Dalili’s work on IP fragmentation [18] explore these protocol-level evasion strategies. Also, SYMTCP [19] focuses on evading deep packet inspection. Furthermore, Handley et al. [20] inspects the evasion of network intrusion detection systems, a concept closely related to bypassing WAFs at the protocol level.
Recent work by Wang et al. [21] explores WAF bypasses through protocol-level evasion, whereas our study takes a distinct approach by targeting content-type-specific parsing discrepancies across headers and bodies. While both use fuzzing, our methodology systematically covers a wider range of content types, leading to the discovery of new bypass classes. We also propose a more detailed classification system with 24 attack types (versus 3 in prior work), validate our techniques on 100 real-world websites beyond controlled environments, and introduce a proof-of-concept normalizer to mitigate these issues. None of which are addressed in Wang et al.’s work.
@@
Payload level bypassing techniques involve transforming the original payload to render it undetectable by the WAF. Obfuscation techniques, such as encoding or encryption, are frequently employed to conceal malicious code from WAF scrutiny. Additionally, evasion techniques, such as altering the order of characters or incorporating whitespace, can be used to confuse the WAF, and facilitate bypassing. Payload size manipulation, such as oversized POST requests that exploited a vulnerability in Google Cloud Armor [22], illustrates how payload modifications can defeat WAF protections. Not all bypasses are feasible using by targeting a single level approach above. Some works involve combining multiple techniques to achieve a bypass. AutoSpear [23] exemplifies a tool that automates WAF bypassing using a combination of architectural, protocol, and payload-level techniques.
@@
4 WAFFLED Design and Methodology
Figure 2 provides an overview of the WAFFLED approach for finding discrepancy-based bypass vectors. The methodology involves generating, mutating, and testing HTTP requests across the WAFs and web application frameworks to identify successful bypasses.
4.1 Input Generation and Mutation
The fuzzer generates valid HTTP requests using a predefined grammar that contains a web attack payload (e.g., SQL injection, Cross-Site Scripting). Mutations are applied everywhere except the attack payload itself, ensuring that bypass attempts focus on discrepancies rather than payload obfuscation. The highlighted part in Listing 3 shows an XSS attack payload, which remains unmodified during mutations. These mutations are expected to give us new discrepancy-based bypass instances. The goal is to confuse WAFs so they fail to properly parse the mutated requests and allow them through, while the web application frameworks, despite the mutations, correctly interpret and execute the attack payload.
4.2 Identifying Successful Bypass Instances
@@ 4.3 Motivating Example
Listing 3: A malicious HTTP request using multipart/form-data with an XSS attack payload.
@@
5boundary=fake-boundary;boundary*0=real-; boundary*1=
6boundary
7
8--fake-boundary
9Content-Disposition: form-data; name="field1"
10
11value1
12--fake-boundary--
13--real-boundary
14Content-Disposition: form-data: name="id"
15
16<script>alert(document.cookie)</script>
17--real-boundary--
@@
5 Experiment Setup
5.1 Tools and Infrastructure for Experiments
We used a modified version of the T-Reqs [17] fuzzer for the experiments. The modifications include adding support for encrypted HTTP requests, and setting content-length dynamically based on the length of the body. We also modified the T-Reqs source code to avoid using < and > characters to parse the grammar, since these characters were also used in the XSS attack payload that was present in generated requests.
@@ 5.2 WAF Configuration and Setup
TABLE I: List of Tested WAFs and Their Supported Request Content-Types (As of July 2024).
@@
Table I lists the WAFs and their supported content-types in our study.
@@
5.3 Web Application Frameworks Setup